OpenStudy (tdskitz):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (tdskitz):

I could really use someone's Help here :/

OpenStudy (google):

 -3(2x + 6) = -12   2x + 6 = 4   2x = -2, x = -1

OpenStudy (google):

 -3(-2x - 6) = -12   -2x - 6 = 4  -2x = 10 ; x = -5

OpenStudy (google):

X= -1 X = -5

OpenStudy (google):

What letter is this?

OpenStudy (google):

fan me if you would like more help i'll always be around

OpenStudy (tdskitz):

Thank you for the help @Google

OpenStudy (google):

No problem any time